OpenStudy (anonymous):

Which shows the factors of x^3-3x^2+4x-12? A. (x^2+4)(x-3) B. (x^2-4)(x+3) C. (x^2-2)(x+6) D. (x^2+2)(x-6)

OpenStudy (cwrw238):

expand each one eg B = x^2*x + 3x^2 - 4x - 12 = x^3 + 3x^2 - 4x - 12 -so its not this one can you continue with this?

OpenStudy (cwrw238):

try expanding A

OpenStudy (anonymous):

X^3-12

OpenStudy (anonymous):

X^3+4x-3x^2-12?

OpenStudy (cwrw238):

ok - you've multiplied x by x^2 and + 4 by -3 but how about x^2 by -3 and + 4 by x?

OpenStudy (cwrw238):

right so is it A?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

That was so easy. Thanks!

OpenStudy (cwrw238):

yea - easy when yu now how. yw